# Optimal energy storage in the Tavis-Cummings quantum battery

### v1

@@ Abstract
The Tavis-Cummings (TC) model, which serves as a natural physical realization of a quantum battery, comprises $N_{b}$ atoms as battery cells that collectively interact with a shared photon field, functioning as the charger, initially containing $n_{0}$ photons. In this study, we introduce the invariant subspace method to effectively represent the quantum dynamics of the TC battery. Our findings indicate that in the limiting case of $n_{0}\!\gg\!N_{b}$ or $N_{b}\!\gg\!n_{0}$, a distinct SU(2) symmetry emerges in the dynamics, thereby ensuring the realization of optimal energy storage. We also establish a negative relationship between the battery-charger entanglement and the energy storage capacity. As a result, we demonstrate that the asymptotically optimal energy storage can be achieved in the scenario where $N_{b}\!=\!n_{0}\!\gg\!1$. Our approach not only enhances our comprehension of the algebraic structure inherent in the TC model but also contributes to the broader theoretical framework of quantum batteries. Furthermore, it provides crucial insights into the relation between energy transfer and quantum correlations.
@@ I Introduction
One notable characteristic of QBs is the phenomenon known as “charging speedup” Hovhannisyan et al. (2013); Binder et al. (2015); Campaioli et al. (2017). When connecting a group of $N_{b}$ QB cells to a shared charger, the charging rate of the QB can potentially scale up to $N_{b}^{2}$yhm et al. (2022), indicating a significant quantum advantage over classical batteries, whose charging speed scales linearly with $N_{b}$. This quantum advantage has been verified in various models, including Dicke QBs Ferraro et al. (2018); Dou et al. (2022a); Zhang and Blaauboer (2023), Sachdev-Ye-Kitaev QBs Rossini et al. (2020), spin-chain QBs Le et al. (2018), cavity spin-chain QBs Dou et al. (2022b) and central-spin QBs Peng et al. (2021). Nevertheless, it is imperative to emphasize that a charging speedup does not inherently ensure the optimal energy storage. Here, the optimal energy storage denotes the complete transfer of input energy from the charger to the battery. To enrich the theory of QBs, we will shift our focus towards an important cavity quantum electrodynamics (cavity-QED) system, the Tavis-Cummings (TC) model, to address the challenge of achieving optimal energy storage.
@@
Figure 1: Schematic illustration of two charging schemes for the Tavis-Cummings battery. (a) Collective charging scheme: $N_{b}$ battery cells (atoms) are commonly interacted with a charger (a cavity photon field). (b) Parallel charging scheme: $N$ single-atom TC models, known as the Jaynes-Cummings model, are placed in parallel. At $t<0$, the battery is in the ground state with the lowest energy. From $0<t<\tau$, the interaction Hamiltonian $H_{I}$ is activated to charge the battery. Finally, at the charging time $\tau$, the interaction is turned off and the energy is stored in the battery.
@@
The TC model serves as a foundational framework for investigating the collective behavior of atoms or qubits in the presence of electromagnetic fields Tavis and Cummings (1968); TAVIS and CUMMINGS (1969); Hepp and Lieb (1973a, b). It constitutes a theoretical underpinning for the fields of quantum optics and cavity-QED Blais et al. (2021); Zare Rameshti et al. (2022); Garraway (2011). Furthermore, it enhances our comprehension of the qubit manipulation and the generation of entangled states, both of which are pivotal components in the advancement of quantum computing and quantum information processing technologies Tessier et al. (2003); Retzker et al. (2007); Fink et al. (2009); López et al. (2007); Steffen et al. (2013); Wang et al. (2020); Leek et al. (2010); Higgins et al. (2014). Moreover, the TC model naturally lends itself to describing a QB, wherein each of the $N_{b}$ atoms serves as an individual battery cell and the optical field serves as the charger, as illustrated in Fig. 1(a). The eigenproblem associated with the TC model can be exactly solved using the algebraic Bethe ansatz Bogolyubov (2000). However, this approach does not alleviate the inherent challenges,exponential wall Kohn (1999) , associated with calculating quantum dynamics, which is at the core of investigating the energy storage problem. Transitioning from the eigenproblem to quantum dynamics is inherently challenging, not to mention the complexity of solving the Bethe ansatz equations. For instance, in Ref. Lu et al. (2021) employing the Bethe ansatz approach, numerical calculations were conducted for a scenario involving 10 QB cells. The study concluded that Fock state is the optimal initial charger state for the maximum stored energy capacity. Reference Andolina et al. (2019) utilized the PYTHON toolbox QuTiP2 to numerically investigate cases in the TC battery where the initial photon number equals the number of battery cells. Their findings revealed that a polynomial increase in the support of the battery state guarantees that the stored work can be completely extracted. For a fixed number of QB cells, Ref. Shi et al. (2022) established a general inverse relationship between the battery-charger entanglement and the extracted work in incoherent QBs. Despite these productive findings, there remains a need for a deeper understanding of the optimal energy storage problem in the TC battery: 1) What conditions must be met for a Fock state to achieve optimal energy storage? 2) How should we address scenarios where the number of battery cells is significantly greater than the initial photon number? 3)Does the negative relation between the entanglement and the energy storage capacity hold as the number of QB cells varies?
@@
In this paper, we address the aforementioned questions for the TC battery. In Sec. II, we begin by introducing the TC model and proceed to formulate the energy storage problem mathematically. Subsequently, we introduce the invariant subspace method to overcome the exponential wall obstruction, allowing us to deal with the dynamics of the TC battery when the charger is chosen in a Fock state. In Sec. III, we first apply the invariant subspace method to analyze the parallel charging scheme. Subsequently, we delve into three collective charging scenarios: $n_{0}\!>\!N_{b}$, $N_{b}\!>\!n_{0}$, and $N_{b}\!=\!n_{0}$. Our analysis rigorously establishes that the emergence of SU(2) symmetry in the limit of $n_{0}\!\gg\!N_{b}$ or $N_{b}\!\gg\!n_{0}$ guarantees the realization of optimal energy storage. Additionally, we provide a verification of the negative correlation between the entanglement and the energy storage capacity. Furthermore, we present numerical evidence that, in the limit of $N_{b}\!=\!n_{0}\!\gg\!1$, the asymptotically optimal energy storage can be attained. In Sect. IV, we provide a discussion regarding the experimental implementation of optimal energy storage within the TC battery and make a conclusion.
@@ II Tavis-Cummings battery
The TC battery consists of $N_{b}$ battery cells, represented by the total spin operators $S^{\alpha}\!=\!\sum_{j=1}^{N_{b}}\sigma_{j}^{\alpha}/2$ $(\alpha\!=\!x,y,z)$, and a charger described by bosonic operators $a$ and $\hat{a}$. The TC Hamiltonian is given by:
@@
The QB is initially prepared in a product state $|\psi(0)\rangle\!=\!|0\rangle_{b}|n_{0}\rangle_{c}$, where the battery is in the ground state of $H_{b}$, i.e., $|0\rangle_{b}\equiv|\downarrow,\ldots,\downarrow\rangle$ and the charger is in a Fock state $|n_{0}\rangle_{c}$. At time $t>0$, the interaction term $H_{I}$ is activated to facilitate the transfer of energy from the charger to the battery. Over a time interval $[0,\tau]$, our objective is to maximize the amount of energy stored in the battery. This optimal time duration $\tau$ is referred to as the “charging time”.
A pivotal step in our analysis involves the calculation of quantum dynamics for the battery state. Due to the U(1) symmetry $[H,S^{z}\!+\!a^{\dagger}a]\!=\!0$ and the local SU(2) symmetries $[H,S^{\alpha}]\!=\!0$ $(\alpha\!=\!x,y,z)$, we observe that the TC battery dynamics can be effectively reformulated within the following subspace:
@@
Within the basis (2), we can express the Hamiltonian (1) as a $(d+1)\times(d+1)$ matrix details are available in the Appendix A):
@@
where $u_{j}\!=\!g\sqrt{j(N_{b}-j+1)(n_{0}-j+1)}$. Assuming that $\bm{H}$ is diagonalized by a unitary matrix $\bm{U}$, i.e., $\bm{H}=\bm{U}\bm{E}\bm{U}^{\dagger}$ where $\bm{E}$ is a diagonal matrix, the matrix representation of the wavefunction for the entire system at time $t$ is given by:
@@
I) The battery absorbs all the energy initially stored in the charger, resulting in zero charger energy after the charging process.
II) The battery absorbs some energy from the charger and reaches a fully charged state $|\uparrow,\ldots,\uparrow\rangle_{b}$.
@@ III.1 Parallel charging: $N_{b}=1$
The parallel charging scheme involves the assembly of $N$ Jaynes-Cummings (JC) models, each of which contains only a single battery cell, as depicted in Fig. 1(b). The JC model is a special case of the TC model by taking $N_{b}\!=\!1$. In this scenario, a straightforward calculation yields the following evolved quantum state (II):
@@
It thus follows from Eq. (12) that the energy stored in battery is given by
@@ III.2 Collective charging: $n_{0}\!>\!N_{b}$
$\bm{U}\!=\!\frac{1}{2\sqrt{2n_{0}-1}}\begin{pmatrix}2\sqrt{n_{0}-1}&\sqrt{2n_{0}}&\sqrt{2n_{0}}\\ 0&\sqrt{4n_{0}-2}&-\sqrt{4n_{0}-2}\\ -2\sqrt{n_{0}}&\sqrt{2(n_{0}-1)}&\sqrt{2(n_{0}-1)}\end{pmatrix},$ (22)
@@
where $h(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$ represents the binary Shannon entropy function. Since $h(x)$ is an increasing function when $20\leq x\leq 1/$ 2 and $n_{0}>N_{b}=2$, we can conclude that entanglement decreases as $n_{0}$ increases. Therefore, we validate the negative relation between the energy storage capacity and the entanglement, see Fig. 2.
@@
A natural question that arises is whether optimal energy storage can be achieved under the condition $n_{0}\!\gg\!N_{b}$, not limited to the specific case of $N_{b}\!=\!2$ as discussed earlier. It appears challenging to address this question by directly solving the eigenproblem of Eq. (8) for arbitrary values of $N_{b}$. Fortunately, we observe that under the condition $n_{0}\!\gg\!N_{b}$ the matrix element $u_{j}\!=\!g\sqrt{j(N_{b}-j+1)(n_{0}-j+1)}$ in Eq. (8) can be approximated as $u_{j}\!\simeq\!g\sqrt{n_{0}-N_{b}/2+1/2}\sqrt{j(N_{b}-j+1)}$. With this approximation, the Hamiltonian (8) can be mapped to the spin-$N_{b}/2$ operator, see the Appendix B:
@@
Now, it is evident that our question has been reduced to a Lie algebra problem, thanks to the emergence of SU(2) symmetry in the limit $n_{0}\!\gg\!N_{b}$. Using the Baker-Campbell-Hausdorff formula, we have
@@
which is extremely accurate even in the finite case $N_{b}\!=\!n_{0}/2\!=\!5$, see Fig 3. Equation (32) indicates that optimal energy storage $\Delta E=N_{b}\omega$ can be achieved at $\tau=\pi/\Omega$ when $n_{0}\!\gg\!N_{b}$.
@@ III.3 Collective charging: $N_{b}>n_{0}$
In Sec. III.2, we employed the invariant subspace method, as given in Eq. (8), to reveal the fundamental algebraic structure that characterizes the quantum dynamical behavior of the TC model. This approach was entirely applicable to the case where $N_{b}>n_{0}$. To delve deeper into the core of our investigation, we will now proceed directly to explore the scenario where $N_{b}\gg n_{0}$ and claim that the optimal energy storage can be realized.
For $N_{b}\!\gg\!n_{0}$, we can approximate $u_{j}$ as $u_{j}\!\simeq\!g\sqrt{N_{b}-n_{0}/2+1/2}\sqrt{j(n_{0}-j+1)}$ and then the Hamiltonian (8) can be approximated as a spin-$n_{0}/2$ operator $\bm{\tilde{J}}^{\alpha}$, i.e.,
@@
which suggests that the optimal energy storage $\Delta E=n_{0}\omega$ can be realized at charing time $\tau=\pi/\tilde{\Omega}$ in the case of $N_{b}\gg n_{0}$. Figure 3 shows the consistency of the limit result with the numerical result obtained also in the finite case $n_{0}\!=\!N_{b}/2\!=\!5$.
@@ III.4 Collective charging: $n_{0}=N_{b}$
The most economical charging scheme is preparing the initial charger with the same energy as the energy storage capacity of the battery, i.e., $n_{0}\!=\!N_{b}$. As we can see in the case of $n_{0}\!\geq\!N_{b}\!=\!2$ (Fig. 2), however, the energy storage capacity is weakest for the $n_{0}\!=\!N_{b}\!=\!2$ case. Hence, we inquire whether, in the limiting case of $n_{0}=N_{b}\gg 1$, we can approach the optimal energy storage. Within this parameter range, no evident approximations for Eq. (8) can be employed to obtain analytical results. Nevertheless, we can also perform an effective numerical calculation based on Eq. (8) to analyze the charging behavior in the case of $n_{0}\!=\!N_{b}$.
@@
Figure 4 illustrates the average entanglement $s(\tau)\!\equiv\!S(\tau)/N_{b}$ and the average stored energy $\Delta\varepsilon(\tau)\!\equiv\!\Delta E(\tau)/N_{b}$ at the charging time $\tau$ for the scenario where $N_{b}\!=\!n_{0}$. The motivation for averaging these quantities stems from the fact that the support of the reduced density matrix, Eq. (11), is determined by $\min\{n_{0},N_{b}\}+1$, which equals $(N_{b}+1)$ in this scenario. Therefore, average quantities enable us to compare cases with varying values of $N_{b}$. Figure 4 shows that the average energy storage capacity $\Delta\varepsilon(\tau)$ initially decreases and then increases with an increase in the size of the QB. Conversely, the average entanglement between the battery and the charger displays the opposite behavior compared to $\Delta\varepsilon(\tau)$. We have previously established a negative relation between the energy storage and the entanglement for a fixed QB size in Ref. Shi et al. (2022). Figure 4 demonstrates that this negative relation persists even when the QB size is not fixed. It is important to emphasize that as $N_{b}$ becomes sufficiently large $(>100)$, the asymptotically optimal energy transfer $(\Delta\varepsilon(\tau)>0.95)$ can be achieved, as depicted in Fig. 4. It is worth noting that a specific perturbation theory, based on polynomially deformed $su$(2) algebras Vadeiko et al. (2003), can be applied to address the eigenproblem of the TC model in the case of $N_{b}=n_{0}$. Exploring the potential applicability of this method to address QBs problems is an interesting topic for further discussion

### v3

@@ Abstract
The Tavis-Cummings (TC) model, which serves as a natural physical realization of a quantum battery, comprises $N_{b}$ atoms as battery cells that collectively interact with a shared photon field, functioning as the charger, initially containing $n_{0}$ photons. In this paper, we introduce the invariant subspace method to effectively represent the quantum dynamics of the TC battery. Our findings indicate that in the limiting case of $n_{0}\!\gg\!N_{b}$ or $N_{b}\!\gg\!n_{0}$, a distinct SU(2) symmetry emerges in the dynamics, thereby ensuring the realization of optimal energy storage. We also establish a negative relationship between the battery-charger entanglement and the energy storage capacity. As a result, we demonstrate that asymptotically optimal energy storage can be achieved in the scenario where $N_{b}\!=\!n_{0}\!\gg\!1$. Our approach not only enhances our comprehension of the algebraic structure inherent in the TC model but also contributes to the broader theoretical framework of quantum batteries. Furthermore, it provides crucial insights into the relation between energy transfer and quantum correlations.
@@ I Introduction
One notable characteristic of QBs is the phenomenon known as “charging speedup” Hovhannisyan et al. (2013); Binder et al. (2015); Campaioli et al. (2017). When connecting a group of $N_{b}$ QB cells to a shared charger, the charging rate of the QB can potentially scale up to $N_{b}^{2}$yhm et al. (2022), indicating a significant quantum advantage over classical batteries, whose charging speed scales linearly with $N_{b}$. This quantum advantage has been verified in various models, including Dicke QBs Ferraro et al. (2018); Dou et al. (2022a); Zhang and Blaauboer (2023), Sachdev-Ye-Kitaev QBs Rossini et al. (2020), spin-chain QBs Le et al. (2018), cavity spin-chain QBs Dou et al. (2022b), and central-spin QBs Peng et al. (2021). Nevertheless, it is imperative to emphasize that a charging speedup does not inherently ensure optimal energy storage. Here, optimal energy storage denotes the complete transfer of input energy from the charger to the battery. To enrich the theory of QBs, we will shift our focus towards an important cavity quantum electrodynamics (cavity-QED) system, the Tavis-Cummings (TC) model, to address the challenge of achieving optimal energy storage.
@@
Figure 1: Schematic illustration of two charging schemes for the Tavis-Cummings battery. (a) Collective charging scheme: $N_{b}$ battery cells (atoms) commonly interact with a charger (a cavity photon field). (b) Parallel charging scheme: $N$ single-atom TC models, known as the Jaynes-Cummings model, are placed in parallel. At $t<0$, the battery is in the ground state with the lowest energy. From $0<t<\tau$, the interaction Hamiltonian $H_{I}$ is activated to charge the battery. Finally, at the charging time $\tau$, the interaction is turned off and the energy is stored in the battery.
@@
The TC model serves as a foundational framework for investigating the collective behavior of atoms or qubits in the presence of electromagnetic fields Tavis and Cummings (1968); TAVIS and CUMMINGS (1969); Hepp and Lieb (1973a, b). It constitutes a theoretical underpinning for the fields of quantum optics and cavity QED Blais et al. (2021); Zare Rameshti et al. (2022); Garraway (2011). Furthermore, it enhances our comprehension of qubit manipulation and the generation of entangled states, both of which are pivotal components in the advancement of quantum computing and quantum information processing technologies Tessier et al. (2003); Retzker et al. (2007); Fink et al. (2009); López et al. (2007); Steffen et al. (2013); Wang et al. (2020); Leek et al. (2010); Higgins et al. (2014). Moreover, the TC model naturally lends itself to describing a QB, wherein each of the $N_{b}$ atoms serves as an individual battery cell and the optical field serves as the charger, as illustrated in Fig.  1(a). The eigenproblem associated with the TC model can be exactly solved using the algebraic Bethe ansatz Bogolyubov (2000). However, this approach does not alleviate the inherent challenges, the exponential wall Kohn (1999) , associated with calculating quantum dynamics, which is at the core of investigating the energy storage problem. Transitioning from the eigenproblem to quantum dynamics is inherently challenging, not to mention the complexity of solving the Bethe ansatz equations. For instance, in Ref. Lu et al. (2021) employing the Bethe ansatz approach, numerical calculations were conducted for a scenario involving ten QB cells. The study concluded that the Fock state is the optimal initial charger state for the maximum stored energy capacity. Reference Andolina et al. (2019) utilized the PYTHON toolbox QUTIP2 to numerically investigate cases in the TC battery where the initial photon number equals the number of battery cells. Their findings revealed that a polynomial increase in the support of the battery state guarantees that the stored work can be completely extracted. For a fixed number of QB cells, Ref. Shi et al. (2022) established a general inverse relationship between the battery-charger entanglement and the extracted work in incoherent QBs. Despite these productive findings, there remains a need for a deeper understanding of the optimal energy storage problem in the TC battery: (1) What conditions must be met for a Fock state to achieve optimal energy storage? (2) How should we address scenarios where the number of battery cells is significantly greater than the initial photon number? (3) Does the negative relation between the entanglement and the energy storage capacity hold as the number of QB cells varies?
@@
In this paper, we address the aforementioned questions for the TC battery. In Sec. II, we begin by introducing the TC model and proceed to formulate the energy storage problem mathematically. Subsequently, we introduce the invariant subspace method to overcome the exponential wall obstruction, allowing us to deal with the dynamics of the TC battery when the charger is chosen in a Fock state. In Sec. III, we first apply the invariant subspace method to analyze the parallel charging scheme. Subsequently, we delve into three collective charging scenarios: $n_{0}\!>\!N_{b}$, $N_{b}\!>\!n_{0}$, and $N_{b}\!=\!n_{0}$. Our analysis rigorously establishes that the emergence of SU(2) symmetry in the limit of $n_{0}\!\gg\!N_{b}$ or $N_{b}\!\gg\!n_{0}$ guarantees the realization of optimal energy storage. Additionally, we provide a verification of the negative correlation between the entanglement and the energy storage capacity. Furthermore, we present numerical evidence that, in the limit of $N_{b}\!=\!n_{0}\!\gg\!1$, asymptotically optimal energy storage can be attained. In Sec. IV, we provide a discussion regarding the experimental implementation of optimal energy storage within the TC battery and make a conclusion.
@@ II Tavis-Cummings battery
The TC battery consists of $N_{b}$ battery cells, represented by the total spin operators $S^{\alpha}\!=\!\sum_{j=1}^{N_{b}}\sigma_{j}^{\alpha}/2$ $(\alpha\!=\!x,y,z)$, and a charger described by bosonic operators $a$ and $\hat{a}$. The TC Hamiltonian is given by
@@
The QB is initially prepared in a product state $|\psi(0)\rangle\!=\!|0\rangle_{b}|n_{0}\rangle_{c}$, where the battery is in the ground state of $H_{b}$, i.e., $|0\rangle_{b}\equiv|\downarrow,\ldots,\downarrow\rangle$ and the charger is in a Fock state $|n_{0}\rangle_{c}$. At time $t>0$, the interaction term $H_{I}$ is activated to facilitate the transfer of energy from the charger to the battery. Over a time interval $[0,\tau]$, our objective is to maximize the amount of energy stored in the battery. This optimal time duration $\tau$ is referred to as the “charging time.”
A pivotal step in our analysis involves the calculation of quantum dynamics for the battery state. Due to the U(1) symmetry $[H,S^{z}\!+\!a^{\dagger}a]\!=\!0$ and the local spin-sector symmetry $[H,\vec{S}^{2}]\!=\!0$, we observe that the TC battery dynamics can be effectively reformulated within the following subspace,
@@
Within the basis (2), we can express the Hamiltonian (1) as a $(d+1)\times(d+1)$ matrix (details are available in Appendix A),
@@
where $u_{j}\!=\!g\sqrt{j(N_{b}-j+1)(n_{0}-j+1)}$. Assuming that $\bm{H}$ is diagonalized by a unitary matrix $\bm{U}$, i.e., $\bm{H}=\bm{U}\bm{E}\bm{U}^{\dagger}$ where $\bm{E}$ is a diagonal matrix, the matrix representation of the wave function for the entire system at time $t$ is given by
@@
(I) The battery absorbs all the energy initially stored in the charger, resulting in zero charger energy after the charging process.
(II) The battery absorbs some energy from the charger and reaches a fully charged state $|\uparrow,\ldots,\uparrow\rangle_{b}$.
@@ III.1 Parallel charging: $N_{b}=1$
The parallel charging scheme involves the assembly of $N$ Jaynes-Cummings (JC) models, each of which contains only a single battery cell, as depicted in Fig.  1(b). The JC model is a special case of the TC model by taking $N_{b}\!=\!1$. In this scenario, a straightforward calculation yields the following evolved quantum state  (II):
@@
It thus follows from Eq. (7) that the energy stored in the battery is given by
@@ III.2 Collective charging: $n_{0}\!>\!N_{b}$
$\bm{U}\!=\!\frac{1}{2\sqrt{2n_{0}-1}}\begin{pmatrix}2\sqrt{n_{0}-1}&\sqrt{2n_{0}}&\sqrt{2n_{0}}\\ 0&\sqrt{4n_{0}-2}&-\sqrt{4n_{0}-2}\\ -2\sqrt{n_{0}}&\sqrt{2(n_{0}-1)}&\sqrt{2(n_{0}-1)}\end{pmatrix},$
@@
where $h(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$ represents the binary Shannon entropy function. Since $h(x)$ is an increasing function when $0\leq x\leq 1/2$ and $n_{0}>N_{b}=2$, we can conclude that entanglement decreases as $n_{0}$ increases. Therefore, we validate the negative relation between the energy storage capacity and the entanglement (see Fig.  2).
@@
[…]
@@
Now, it is evident that our question has been reduced to a Lie algebra problem, owing to the emergence of SU(2) symmetry in the limit $n_{0}\!\gg\!N_{b}$. Using the Baker-Campbell-Hausdorff formula, we have
@@
which is extremely accurate even in the finite case $N_{b}\!=\!n_{0}/2\!=\!5$ (see Fig.  3). Equation (20) indicates that optimal energy storage $\Delta E=N_{b}\omega$ can be achieved at $\tau=\pi/\Omega$ when $n_{0}\!\gg\!N_{b}$.
@@ III.3 Collective charging: $N_{b}>n_{0}$
In Sec.  III.2, we employed the invariant subspace method, as given in Eq.  (3), to reveal the fundamental algebraic structure that characterizes the quantum dynamical behavior of the TC model. This approach was entirely applicable to the case where $N_{b}>n_{0}$. To delve deeper into the core of our investigation, we will now proceed directly to explore the scenario where $N_{b}\gg n_{0}$ and claim that optimal energy storage can be realized.
[…]
@@
which suggests that optimal energy storage $\Delta E=n_{0}\omega$ can be realized at charging time $\tau=\pi/\tilde{\Omega}$ in the case of $N_{b}\gg n_{0}$. Figure  3 shows the consistency of the limit result with the numerical result obtained also in the finite case $n_{0}\!=\!N_{b}/2\!=\!5$.
@@ III.4 Collective charging: $n_{0}=N_{b}$
The most economical charging scheme is preparing the initial charger with the same energy as the energy storage capacity of the battery, i.e., $n_{0}\!=\!N_{b}$. As we can see in the case of $n_{0}\!\geq\!N_{b}\!=\!2$ (Fig.  2), however, the energy storage capacity is weakest for the $n_{0}\!=\!N_{b}\!=\!2$ case. Hence, we inquire whether, in the limiting case of $n_{0}=N_{b}\gg 1$, we can approach optimal energy storage. Within this parameter range, no evident approximations for Eq.  (3) can be employed to obtain analytical results. Nevertheless, we can also perform an effective numerical calculation based on Eq.  (3) to analyze the charging behavior in the case of $n_{0}\!=\!N_{b}$.
@@
Figure  4 illustrates the average entanglement $s(\tau)\!\equiv\!S(\tau)/N_{b}$ and the average stored energy $\Delta\varepsilon(\tau)\!\equiv\!\Delta E(\tau)/N_{b}$ at the charging time $\tau$ for the scenario where $N_{b}\!=\!n_{0}$. The motivation for averaging these quantities stems from the fact that the support of the reduced density matrix, Eq.  (6), is determined by $\min\{n_{0},N_{b}\}+1$, which equals $(N_{b}+1)$ in this scenario. Therefore, average quantities enable us to compare cases with varying values of $N_{b}$. Figure  4 shows that the average energy storage capacity $\Delta\varepsilon(\tau)$ initially decreases and then increases with an increase in the size of the QB. Conversely, the average entanglement between the battery and the charger displays opposite behavior compared to $\Delta\varepsilon(\tau)$. We have previously established a negative relation between the energy storage and the entanglement for a fixed QB size in Ref.  Shi et al. (2022). Figure  4 demonstrates that this negative relation persists even when the QB size is not fixed. It is important to emphasize that as $N_{b}$ becomes sufficiently large (>100)(>100)( > 100 ), the asymptotically optimal energy transfer $[\Delta\varepsilon(\tau)>0.95]$ can be achieved, as depicted in Fig.  4. It is worth noting that a specific perturbation theory, based on polynomially deformed su(2) algebras Vadeiko et al. (2003), can be applied to address the eigenproblem of the TC model in the case of $N_{b}=n_{0}$. Exploring the potential applicability of this method to address QB problems is an interesting topic for further discussion.